Q . THE EQUATIONS OF THE LINES REPRESENTING THE SIDES OF A TRIANGLE ARE-
3x – 4y = 0 , x + y =0  AND 2x – 3y =7. THE LINE- 3x+ 2y =0 ALWAYS PASSES THROUGH THE – 
  1. INCENTRE
  2. CENTROID
  3. CIRCUMCENTRE
  4. ORTHOCENTRE

Grade:Select Grade

1 Answers

vardhan
9 years ago
 The lines 3x – 4y  =0 and x+y=0 intersect each other at    (0,0). 
Also the line 3x + 2y = 0 passes through (0,0) and is perpendicular to the third line 2x – 3y=7 . 
Therefore it passes through ORTHOCENTRE as it is perpendicular to third line and passes through the point of intersection of first and second line.